OpenStudy (anonymous):

Write the formula given the following arithmetic sequences. 4, 8, 12, 16,… Arithmetic sequences are not my thing.... -.-

OpenStudy (anonymous):

i guess your looking for 4

OpenStudy (anonymous):

well i need a forumla so i dont think 4 will do... List of answers: tn = 2 – 2(n – 1) tn = 2 + 4(n – 1) tn = 4 + 4(n – 1) tn = 4 – 4(n – 1) Please explain before answering!

OpenStudy (anonymous):

tn=4+4(n-1) would be my guess

OpenStudy (anonymous):

mmkay i'll trust you on this one lol :)

OpenStudy (anonymous):

Good luck